OpenStudy (anonymous):

What is the slope of a line parallel to the line whose equation is 2x – 3y = 6? A. 1/5 B. 2/3 C. 4/3 D. 3/3

OpenStudy (anonymous):

just solve for y to get it in y=mx+b form and m will be answer

OpenStudy (anonymous):

ummm what?

OpenStudy (anonymous):

2x – 3y = 6 solve for y – 3y = 6- 2x subtract 2x from both sides y=-6/3 - (-2x/3)

OpenStudy (anonymous):

y=-6/3 - (-2x/3) y=-2 + 2/3 x ^ the - and - cancel out to make + y=2/3x-2 rearrange y= m x + b answer is 2/3

OpenStudy (anonymous):

ohhhhh! u think you can help me with one more?

OpenStudy (anonymous):

ok just one more

OpenStudy (anonymous):

Write the inequality for the graph. The points are (0,4) and (2,0). A. y < -2x+4 (The symbols sopose to be a less than or equal to sign.) B. y < -2x+4 C. y > -2x+4 (This one is sopose to be a greater than or equal to sign.) D. y > -2x+4

OpenStudy (anonymous):

where is the graph?

OpenStudy (anonymous):

thats not ittt

OpenStudy (anonymous):

^^^^^^^^^^^^^^^^^^^^^^^^that is^^^^^^^^^^^^^^^^^^^^

OpenStudy (anonymous):

Are you still here?

OpenStudy (anonymous):

yes i am working on i

OpenStudy (anonymous):

Okay i put the graph up

OpenStudy (anonymous):

OOK the line is solid so it has to be A or C other wise if it was just < or > it be a dashed line -2x+4 this is your equation 0 pick an easy point for x which is 0 4 is the result 3 Pick a point in the shaded area on the y axis which is also point 0. So I pick 3 4 > 3 compare the result with the point since 4 is y and it greater than any point under its shaded area it is D. y > -2x+4